# Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression

### v2

@@ 4. Evaluation & Analysis
variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
Table 1 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\mathrm{variant}}-\mathrm{Acc}_{\mathrm{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants. All models achieve Acc=1.0000=1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$). By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links. Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
For logic-preserving single-law rewrites (Variant 4), performance is largely preserved: most conditions remain at 1.0000 accuracy, with only small degradations in a few settings (e.g., BERT under commutativity: 0.9925, $\Delta=-0.0075$; Qwen2 under implication-to-disjunction: 0.9525, $\Delta=-0.0475$). Under multi-law equivalence stacking (2–5 laws), robustness becomes model-dependent: BERT matches the base condition (1.0000), TinyLlama shows only a marginal reduction (0.9925, $\Delta=-0.0075$), whereas Qwen2 exhibits a pronounced decrease (0.6450, $\Delta=-0.3550$). This contrast suggests that compositional logical rewrites can introduce effective difficulty beyond single transformations, even when semantic equivalence is preserved.
Generalisation and Sensitivity
Variants 1–3 isolate model sensitivity to perturbations that preserve versus disrupt derivability. In Variant 1 (redundant rule deletion), all models match the base accuracy, consistent with the expectation that removing unused rules should not affect entailment. In Variant 2 (essential rule deletion), accuracy drops to near-chance levels, suggesting limited premise tracking: models often behave as if the missing link were still present and tend to “complete” the original inference pattern even when intermediate support is removed. This behaviour is consistent with reliance on learned chain templates rather than verifying support for each queried conclusion. In Variant 3 (contradictory evidence), all models score 0.0000, indicating that they do not reliably adopt contradiction-aware conservative reasoning under the benchmark’s scoring semantics.
For logic-preserving rewrites (Variant 4), performance is largely preserved for single-law transformations, with only minor degradations in a few cases (e.g., implication-to-disjunction or commutativity). However, multi-law stacking exposes model-dependent robustness: BERT remains at the base level, TinyLlama shows only marginal reduction, whereas Qwen2 degrades substantially. This suggests that composing multiple equivalence transformations introduces an effective distribution shift beyond isolated rewrites, likely through increased operator density and structural complexity despite semantic equivalence.
Implications and Diagnostic Value
The variant-wise profile provides a compact diagnostic of three failure modes: (i) missing-link sensitivity under essential rule deletion, indicating weak verification of minimal supporting premises; (ii) inconsistency handling under contradiction injection, where models fail to revise or withhold conclusions when counter-evidence is present; and (iii) compositional sensitivity, where multi-law rewrites can degrade performance even when each rewrite is semantics-preserving. The strong gap between single-law and multi-law conditions cautions against equating robustness to isolated perturbations with robustness under composition. These diagnostics motivate targeted interventions such as deletion-aware training, explicit proof/rationale verification, and stress-testing robustness under composed transformations rather than single perturbations alone.
Conclusion
This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations of rule-based reasoning. By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolate behaviours that are often obscured in more heterogeneous benchmarks. Across BERT, Qwen2, and LLaMA-like models, we observe Acc=1.0000=1.0000 on the base split and no degradation under redundant rule deletion, but pronounced failures under essential rule deletion and uniform breakdown under explicit contradictions. For equivalence-based rewrites, performance is largely preserved for single-law transformations, while composed (multi-law) rewrites reveal model-dependent sensitivity. These findings underscore the gap between robustness to semantic-preserving reformulations and reliable reasoning under missing, inconsistent, or compositionally transformed evidence, and motivate methods that explicitly address premise tracking and inconsistency management.
@@ Appendix A Appendix
Base Example: Complex Dilemma Reasoning Structure
@@
Variation 1: Rule Reduction with Same Conclusions
@@
Variation 2: Rule Equivalence with Different Conclusions
@@
Variation 3: Rule Interference with Contradictory Conclusions
@@
Reasoning Process and Analysis
This variation introduces contradictory interference, which challenges the AI’s ability to resolve conflicts within the logical structure. The fact “Anne is green or blue,” combined with the original chain of reasoning (Rules 1-6), leads to the conclusions that Anne is cold, rough, young, and nice. However, the fact “Anne is not cold or not nice” introduces a conflict, as it implies that being nice would mean Anne is not cold. This creates an inconsistency that different AI systems might resolve differently:
  • Conservative approach: Detect inconsistency and withhold conclusions
@@
The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for inconsistency detection and resolution.

### v3

@@ 4.1 Main Results: Conflict-Aware Fusion Performance
variant2 (Essential Deletion) 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
variant3 (Contradiction) 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
variant4-multi (Stacked Laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
Table 1: Baseline Performance: Accuracy and deviation ($\Delta$) across structural variants without fusion optimization.
To resolve these failure modes, we propose the Fusion-Conflict framework. As summarized in Table 2, our method sets a new benchmark for logical reasoning, achieving 1.0000 on both the base task and the contradiction stress test (Variant 3).
Table 2: Final Performance: Comparison of All Methods (Accuracy). Fusion-Conflict (Ours) achieves 100% accuracy on contradictions while recovering base performance.
Method Base Acc Var 2 (Rule Removal) Var 3 (Contradiction) Rank
Stage 1 (SFT Baseline) 0.512 0.250 0.210 7
DPO (Direct Alignment) Rafailov et al. (2023) 0.475 0.267 0.510 6
CoT (Standard) Wei et al. (2022) 0.500 0.390 0.865 5
Mixed-Aug (Data Augmentation without CoT) 0.525 0.405 0.972 4
RA-CoT (Standard CoT + DPO) 0.263 0.593 0.690 3
Fusion-LRA (Conflict-Aware SFT) 0.988 0.753 0.705 2
[…]
4.2 Experimental Setup and Hyperparameters
Dataset and Models The benchmark consists of multi-step inference chains with 100 base groups (80 for training, 20 for testing). We evaluate BERT-base, Qwen2-1.5B, and TinyLlama-1.1B. All training uses Low-Rank Adaptation (LoRA) for parameter efficiency.
Stage 1 Training Details For the initial rule-prediction phase, the hyperparameters used for the generative models are detailed in Table 3.
Hyperparameter Value
Epochs 3
Learning Rate $2.0\times 10^{-5}$
Batch Size 4
Max Sequence Length 512
Optimizer AdamW (Mixed Precision)
LoRA Config r: 8, $\alpha$: 16, dropout: 0.05
Table 3: Detailed Stage 1 Training Hyperparameters.
4.3 Ablation and Incremental Optimization
To understand the evolution of our methodology, we track the performance of intermediate strategies. Table 4 highlights the impact of Stage 1 pre-training alone, which shows that while rule prediction helps slightly with Variant 2, it causes significant degradation in base task accuracy without subsequent fusion stages.
Ablation: Effect of Stage 1 Pre-training on Generalisation
BERT Qwen2 TinyLlama
Split Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
base 1.0000 0.0000 0.5250 -0.4750 0.5375 -0.4625
variant2 (Essential Deletion) 0.2500 -0.7500 0.4050 -0.1200 0.5325 -0.0050
variant3 (Contradiction) 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
Table 4: Ablation Results: Stage 1 pre-training moderately improves Variant 2 robustness for generative models but severely degrades base performance.
Finally, Table 5 illustrates the trade-offs encountered during the development of Fusion-LRA. This intermediate step achieved an optimal balance (Rank 1 at the time) by recovering 98.8% base accuracy, providing the foundation for the final Conflict-Aware architecture.
Method Base Acc Var 2 Var 3 Rank
Mixed-Aug 0.525 0.405 0.972 3
RA-CoT 0.263 0.593 0.690 2
Fusion-LRA 0.988 0.753 0.705 1
Table 5: Iterative Comparison: Fusion-LRA vs earlier iterations.
4.4 Is DPO Necessary? (Impact Analysis)
A key question in our optimization pipeline is the specific contribution of Stage 2 Reinforcement Learning (DPO) versus Supervised Fine-Tuning (SFT). Our results provide a definitive answer.
As shown in Table 2, Fusion-LRA (SFT Only) achieved a strong Base Accuracy of 0.988 but plateaued at 0.705 for Contradiction handling (Variant 3). This indicates that while SFT can teach the model the form of the contradiction check, the model still occasionally “hallucinates” a successful check even when a contradiction exists (a phenomenon we term “shallow alignment”).
In contrast, Fusion-Conflict (SFT + DPO) pushes the Variant 3 accuracy to 1.0000. By explicitly penalizing trajectories where the model fails to halt on contradictions (using DPO preference pairs), we sharpen the decision boundary of the Contradiction Detection. Consequently, DPO does not merely incrementally improve performance; it acts as the critical robustness hardener, bridging the final gap from “mostly consistent” to “logically sound.”
5 Discussion and Implications
5.1 Comparative Analysis of Robustness Strategies
Our experimental results, summarized in Table 2, reveal a clear evolutionary path from baseline models to the proposed Conflict-Aware Fusion framework.
Initially, baseline models (Stage 1) demonstrated high task proficiency but were essentially ”logic-blind” to contradictions, scoring near 0.210 on Variant 3. While Mixed-Aug specialized in identifying contradictions (reaching 0.972), it failed to scale to complex reasoning chains. The RA-CoT framework introduced more rigor but suffered from an ”over-conservatism” trap, where the model preferred to output ”False” for nearly everything, causing Base Accuracy to plummet to 0.263.
The Fusion-Conflict approach successfully breaks this trade-off. By achieving 1.0000 accuracy in both Base and Variant 3 (Table 2), it proves that the model no longer relies on opportunistic deduction. Instead, it has internalized the ”Contradiction Detection” as a prerequisite for logic, effectively solving the ”Logic Inertia” problem that plagued all previous iterations.
5.2 The Success of the Cognitive Structure Hypothesis
The decisive performance of Conflict-Aware Fusion provides strong empirical support for our Cognitive Structure Hypothesis. As demonstrated by the comparison between Fusion-LRA and Fusion-Conflict, simply improving data balance (LRA) leads to substantial performance gains (0.705 on V3); however, these gains remain inferior to those achieved by explicitly modeling cognitive conflict.
This suggests that for LLMs, ”Reasoning” is not just a statistical continuation of tokens, but a process that can be steered by structural inductive biases. The ”Step 1: Verify facts” preamble acts as a functional circuit breaker, preventing the model’s System 1 (intuition/pattern matching) from overriding the factual reality of the premises.
5.3 Generalisation and Sensitivity
Variants 1–3 isolate model sensitivity to perturbations that preserve versus disrupt derivability. In Variant 1 (redundant rule deletion), our Fusion model maintains Acc=1.000, confirming that the internal contradiction detection does not introduce fragility to noise.
However, Variant 2 (Essential Rule Removal) remains the most challenging stress test. While Fusion-Conflict achieves a high score of 0.735, it still shows a delta compared to the base set. This indicates that while the model is excellent at detecting explicit contradictions (V3), it is still slightly prone to ”logical leaps” when an inferential link is missing but the overall pattern looks familiar. This diagnostic highlights a future research direction: enhancing the model’s sensitivity to logical gaps as much as it is now sensitive to logical conflicts.
5.4 Implications for Reliable AI Deployment
The transition from RA-CoT’s over-conservatism to Fusion-Conflict’s precision has significant implications for high-stakes deployment:
  1. 1.
    Process over Data: Improving reasoning reliability is better achieved by refining the reasoning process (via structured CoT) than by merely expanding variables in the training set.
  2. 2.
    DPO as a Logical Refiner: Our use of DPO (Stage 2) demonstrates that preference alignment can be used to punish hallucinatory logic, shifting the model from ”answering” to ”verifying.”
  3. 3.
    Diagnostic Value: Our multi-variant benchmark allows researchers to distinguish between a model that ”knows the rules” and a model that ”knows when the rules don’t apply.”
5.5 Real-World Validation: LogicNLI & MNLI
To verify the generalization of our Cognitive Structure Hypothesis, we extended our evaluation to real-world datasets, specifically LogicNLI Tian et al. (2021) (a First-Order Logic NLI benchmark) and the contradiction subset of MNLI Williams et al. (2018). Unlike synthetic rules, these datasets feature diverse linguistic patterns and implicit logical structures.
We fine-tuned a lightweight 0.5B model (Qwen2.5-0.5B) using our Conflict-Aware Template (Stage 1 SFT). The model was taught to explicitly output “Step 1: Detect Contradiction” before predicting entailment or contradiction. Initial results are highly promising: the model successfully learned to invoke the “Halt” mechanism upon detecting conflicts in LogicNLI samples.
Method LogicNLI MNLI (Con)
Standard SFT (Baseline) 58.5% 34.1%
Mixed-Aug (Data Only) 65.2% 41.5%
Fusion-LRA (No DPO) 92.4% 78.3%
[…]
Table 6: Zero-shot transfer performance. “MNLI (Con)” refers to accuracy specifically on the contradiction subset of the Multi-Genre NLI dataset, which tests the model’s ability to reject inconsistent premises in open-domain text.
This demonstrates that the specific structural prior we propose—separating verification from deduction—is not limited to synthetic logic but is a viable strategy for robust natural language inference in complex domains.
6 Conclusion
In this work, we present a controlled framework for analyzing and enhancing the structural robustness of rule-based reasoning in large language models. By applying systematically designed perturbations, we demonstrate that contemporary models—including BERT, Qwen2, and TinyLlama—maintain high accuracy under standard and form-equivalent conditions, yet consistently fail when essential rules are removed or explicit contradictions are introduced. This behavior reflects a persistent reliance on forward inferential continuation rather than explicit premise validation.
To address this limitation, we introduce the Conflict-Aware Fusion framework, which enforces an explicit separation between premise verification and deductive execution within the reasoning trajectory. Empirical results show that this structural intervention substantially improves robustness under rule deletion and fully resolves contradiction-induced failure, achieving 1.0000 accuracy in contradiction handling while preserving base-task performance. These findings provide direct empirical support for the Cognitive Structure Hypothesis: reliable multi-step reasoning depends not only on data exposure but also on the structural organization of the reasoning process.
Beyond these technical results, our work demonstrates that the gap between semantic-form invariance and logical-content reliability can be systematically bridged through architectural priors in Chain-of-Thought prompts. By operationalizing explicit premise verification, Conflict-Aware Fusion offers a scalable blueprint for doubt-aware AI systems capable of disciplined, evidence-driven inference. Overall, this work contributes both a diagnostic benchmark and a constructive methodology for developing LLMs that reason with verifiable logical rigor rather than surface-level pattern completion.
@@ Appendix A Appendix
A.1 Base Example: Complex Dilemma Reasoning Structure
@@
A.2 Variation 1: Rule Reduction with Same Conclusions
@@
A.3 Variation 2: Rule Equivalence with Different Conclusions
@@
A.4 Variation 3: Rule Interference with Contradictory Conclusions
@@
A.5 Reasoning Process and Analysis
This variation introduces contradictory interference, which challenges the AI’s ability to resolve conflicts within the logical structure. The fact “Anne is green or blue,” combined with the original chain of reasoning (Rules 1-6), leads to the conclusions that Anne is cold, rough, young, and nice. However, the fact “Anne is not cold or not nice” introduces a conflict, as it implies that being nice would mean Anne is not cold. This creates an contradiction that different AI systems might resolve differently:
  • Conservative approach: Detect contradiction and withhold conclusions
@@ A.5 Reasoning Process and Analysis
The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for contradiction detection and resolution.
@@ A.6 Comparative Analysis
Split BERT Qwen2 TinyLlama
Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
base 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant1 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
Table 8 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\text{variant}}-\mathrm{Acc}_{\text{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants. All models achieve Acc = 1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$). By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links. Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
Figure 2: An example where all top-tier models failed the Human Last Exam.
Figure 2 shows a screenshot of the AGI Safe AI dashboard under the section “II. AI Answers”, where multiple state-of-the-art language models (including Claude Sonnet 4.5, GPT-4.1, GPT-5.2, Claude Opus 4.5, and Gemini 3 Pro Preview) are evaluated on a logical reasoning question involving a set of premises about “Anne” and four sub-questions (Q1–Q4).
This example 2 evaluates reasoning with a disjunctive color premise and a conflicting constraint. From the fact “Anne is green or blue” and the rules stating that being green or being blue implies being cold, it follows that Anne is cold in either case. Applying the rule “if someone is cold then they are rough,” we can further derive that Anne is rough.
However, the additional fact “Anne is not cold or not nice” introduces a conflict with the derived conclusion that Anne is cold. Given $\mathrm{Cold}(\mathrm{Anne})$, the disjunction $\neg\mathrm{Cold}(\mathrm{Anne})\lor\neg\mathrm{Nice}(\mathrm{Anne})$ forces the conclusion that Anne is not nice. As a result, the statement that Anne is nice is contradicted.
Regarding youth, the rules only support forward implications: being young implies being cold and being nice. There is no rule that allows inferring that Anne is young from coldness, roughness, or any other established fact. Therefore, Anne being young cannot be derived from the given information.
Different reasoning frameworks may handle the inconsistency differently (e.g., conservative systems may block some inferences, whereas paraconsistent systems may continue reasoning in the presence of contradictions). Under standard conservative reasoning, Anne is cold and rough are supported, Anne is young is undetermined, and Anne is nice is contradicted.
The core interaction can be summarized as:
$C_{a}\land(\neg C_{a}\lor\neg N_{a})\vdash\neg N_{a}$
  • $C_{a}=\mathrm{Cold}(\mathrm{Anne})$
  • $N_{a}=\mathrm{Nice}(\mathrm{Anne})$
Each model’s output is displayed in separate panels with its predicted answers and reasoning. The results indicate that the models produce inconsistent and conflicting conclusions, with some models claiming all statements are true while others detect inconsistency in the premises. At the bottom, the evaluation summary reports that 0 out of 5 models answered correctly, suggesting that this example is challenging even for top-tier models and highlights their limitations in formal logical reasoning and consistency checking.